Many chemical and refining companies use "the rule of point six" to estimate the cost of new equipment. According to this rule, if a piece of equipment (such as a storage tank) originally cost C dollars, then the cost of similar equipment that is x times as large will be approximately x * 0.6 * C dollars. For example, if the original equipment cost C dollars, then new equipment with twice the capacity of the old equipment (x = 2) would cost 2 * 0.6 * C = 1.516 * C
dollars—that is, about 1.5 times as much. Therefore, to increase capacity by 100% costs only about 50% more. Use the rule of .6 to find how costs change if a company wants to quadruple (x = 4) its capacity. (Round your answer to one decimal place.)
C dollars

Answer by Theo(13342)

i believe what you are saying is:

if a piece of equipment originally cost c dollars, then the cost of simillar equipment that is x times as large will be approximately x * .6c dollars.

your example states that, if the original equipment cost c dollars, then new equipment with twic3e the capacity of the old equipment would cost 2 * .6 * c dollars = 1.516 * c dollars.

unfortunately, as far as i can tell, 2 * .6 * c dollars is equal to 1.2 * c dollars.

i don't know where 1.516 * c dollars came from.
it is very confusing to me, unless i am able to understand how you arrived at that figure.

if i assume the formula is as you stated, then i would go with 2 * .6 * c 1.2 * c.

if a new equipment is 4 times as large, then the rule of point six would state that the cost of the equipment would be equal to 4 * .6 * c = 2.4 * c dollars.

let me know if you have any questions and how you obtained 1.516 rather than 1.2 when the new equipment was 2 times as large.